How much work must be done to a proton to accelerate it from rest to 0.998c.


1
Expert's answer
2022-05-30T15:17:56-0400
"W=\frac{mc^2}{\sqrt{1-v^2/c^2}}-mc^2=mc^2(\frac{1}{\sqrt{1-v^2/c^2}}-1)"

"W=1.67*10^{-27}*(3*10^8)^2(\frac{1}{\sqrt{1-0.998^2}}-1)\\
=2.2*10^{-9}\:\rm J"